OpenStudy (anonymous):

I'm supposed to use the quadratic formula to solve x^2 - 7x - 6 = 0 Ufortunatly it's a subject I've been having problems with since about halfway through, and now that I'm at the end I'm sorta freaking out.

OpenStudy (anonymous):

x^2 - 7x -6 = 0 Here, a = 1 b = -7 c = -6 Then you put those values in this equation and find the answer. x = [ -b ± sqrt(b^2 - 4ac) ] / 2a

OpenStudy (anonymous):

http://www.purplemath.com/modules/quadform.htm Because I cannot write the equations correctly here for some reason...

OpenStudy (anonymous):

OH, I would have failed horribly, that was actually the only equation I understood very well, I mixed up my names I guess...

OpenStudy (anonymous):

\[x =7+ \sqrt{73}/2 or x =7- \sqrt{?73}/2\]